OpenStudy (anonymous):

Simplify the radical expression. -2√2p * 2√22 A. √44p B. -8√11p C. -4√44p D. -8√11p^2

OpenStudy (anonymous):

Multiply the outside and inside separately. \[-2\sqrt{2p} \times 2\sqrt{22} \implies -2 \times2(\sqrt{2p \times 22}) \implies?\]